OpenStudy (anonymous):

find max or min value of function tell whether this value is a max or min. f(x)=-2x^2-4x+3

OpenStudy (anonymous):

max is at x=-1 use your calculator to graph

OpenStudy (anonymous):

x = -b/2a a = -2 b = -4 c = 3 x = 4/-4 = -1 Plug in (-1) for x and you get 5 Max at (-1,5) it is a max because -2x^2 is negative.

OpenStudy (anonymous):

f(x)=-3(x+3)^2-7 same rules